OpenStudy (anonymous):

2y^2-8y+5=0 <--- solve by the quadratic formula , set the equation to "0" first. must be in simplest radical form, show work please

hero (hero):

If I solve it, it will only confuse you. I have this crazy idea that all quadratics are factorable

OpenStudy (anonymous):

ohh:p

OpenStudy (anonymous):

-8+\[8\pm \sqrt{(-8)^2-4(2)(5)}\] Im doing the numerator first

hero (hero):

You could have at least posted the actual quadratic formula first

OpenStudy (anonymous):

sorry http://en.wikipedia.org/wiki/Quadratic_equation

OpenStudy (anonymous):

you solve the inside of square root you will get \[8\pm \sqrt{64-40}\]

OpenStudy (anonymous):

\[8\pm \sqrt{24}\]

OpenStudy (anonymous):

\[8\pm2\sqrt{6}\]

OpenStudy (anonymous):

now all of this is over the denominator 2(2) correct it simplifies to \[2\pm (\sqrt{6}\div2)\]

OpenStudy (anonymous):

thanks